OpenStudy (anonymous):

Determine which answer choice below is equivalent to the given expression by completing the square. -x^2 -10x + 19

OpenStudy (anonymous):

where are the choices?

OpenStudy (anonymous):

\[\huge (x+5)^2=44\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\huge (x+5)=\pm \sqrt{44}\]